# Lookup multivariate Kolmogorov-Arnold Networks

### v1

@@ 4.3.1 CIFAR-10
Our backbone architecture consists of five $2\times 2$ convolutions, each with stride $2$, and two fully-connected layers at the end. Since the resolution of CIFAR-10 images is $32\times 32$, where $32=2^{5}$, five $2\times 2$ convolutions with stride $2$ transform the spatial dimensions of an image exactly to $1\times 1$. All the layers use the same width (= number of filters in case of convolutions, and hidden dimension in case of fully-connected layers), which we vary for both families of the models. In other aspects, the models are similar to those we employed in previous sections - we use batch normalizations with affine transforms for MLP-CNNs, and without for lmKAN-CNNs; MLP-CNNs use ReLU activations, while lmKAN-CNNs do not require additional activation layers.
The dataset comes with pre-defined full training and test subsets. We split the full training subset into training and validation parts in a 90%/10% ratio. Our augmentation pipeline consists of established techniques, such as RandAugment Cubuk et al. (2020), MixUP Zhang et al. (2017), CutMix Yun et al. (2019), and a few others. Further details about the augmentations and other aspects of our fitting setup are available in appendix. F.4.
Our findings are illustrated in the upper row of Fig. 11. Similarly to previous experiments, lmKAN-based CNNs were found to be more FLOPs efficient compared to classical MLP-based CNNs at the same accuracy level. The observed speed-up factor of 1.6-2.1× is not so dramatic as we report in sections 4.1 and 4.2, but still substantial. We have not yet implemented dedicated CUDA kernels for efficient inference of lmKAN-based convolutions. Any type of convolution can be cast to a fully-connected layer by the corresponding memory manipulations, which we followed during fitting.
@@
Appendix F.4 contains ablation studies on 1) the effect of the number grid intervals $G$ and 2) the effect of the strength of Hessian regularization $\lambda$.
@@ 4.3.2 ImageNet
To limit the computational cost, we downsampled the images to $81\times 81$ ($81=3^{4}$) pixels. Our backbone, illustrated in the right panel of Fig. 10, consists of four convolutional layers with the $3\times 3$ kernel size and stride $3$ and two fully-connected layers. Contrary to the CIFAR-10 experiment, we progressively increase the number of filters as the spatial resolution of the image decreases through the neural network. Our augmentation pipeline consists of the same techniques as the one we use for CIFAR-10, but with different hyperparameters, see more details in Appendix F.5:. Since the test subset is not publicly available we use the validation accuracy as the target metric. The performance of the models is summarized in the bottom row of Fig. 11. The observed efficiency gains of 1.7× are in line with those for CIFAR-10.
@@ 4.4 Comparison with FastKAN
We use the training script [^3] https://github.com/ZiyaoLi/fast-kan/blob/17b65401c252334fffb5e63c9852dd8316d29e69/examples/train_cifar10.py for the CIFAR-10 dataset available in the FastKAN GitHub repository Li (2024b) as the basis for the comparison of lmKAN and FastKAN. However, we provide several modifications to the pipeline.
@@
We extend the script by the same augmentational pipeline as we used in Sec. 4.3 for the CIFAR-10 dataset. We observed that because of augmentations, one has to fit the model for a longer time, so the training budget was substantially increased. Additional changes are detailed in appendix F.6.
Figure 12: Comparison of lmKAN and FastKAN within the fully-connected framework on the CIFAR-10 dataset. Both models were fit with 1k and 10k epochs.
These modifications significantly improve the performance of FastKAN models ($\approx 54-55\%$ validation accuracy in the original script), see Fig. 12. Furthermore, even the MLP baseline of the same shape yields better accuracy compared to the performance of the FastKAN model in the original script.
@@
The primary observation is that the performance of FastKAN models degrades for excessively fine grid resolutions. For the 1k epochs training budget, the final model is even less accurate compared to the MLP baseline. For the 10k epochs, the effect is less pronounced, but it still takes place. For lmKAN models, this degradation is much less severe, if present at all.
The number of parameters per function scales quadratically with grid resolution for lmKANs and linearly for FastKANs. Thus, although the rightmost lmKAN model in Fig. 12 uses more parameters per function, it operates on a much coarser grid — only $G=16$ grid intervals per dimension, compared with num_grids=200num\_grids=200 for FastKANs. We hypothesize that, for a comparable parameter budget per function, this coarser grid enforces a lower-frequency function class and, in turn, contributes to the superior training stability observed for lmKANs.
Another distinct feature of Fig. 12 is that lmKANs achieve notably better accuracy compared to FastKANs, even when the latter are evaluated with a very rich parametrization of inner functions. Additionally, note that the number of trainable functions s $N_{in}/2]N_{out}$ and $N_{in}N_{out}$ in an lmKAN and FastKAN layer, respectively. Thus, the same number of parameters per function corresponds to a two times larger FastKAN model overall.
To sum up, these findings reinforce the intuitional considerations given in Sec. 3 and suggests that building blocks of multivariate trainable functions are indeed more effective.
@@ 6 Summary
High-dimensional linear mappings are the cornerstone of modern deep learning, dominating both the parameter count and the computational cost in most models. We introduce lookup multivariate Kolmogorov–Arnold Networks (lmKANs) as a drop-in replacement that offers a substantially better capacity–inference cost trade-off. Across all experiments, lmKANs were Pareto-optimal in the inference FLOPs–performance plane. The efficiency gains are task-dependent: for general high-dimensional function approximation, modelled as a distillation from a large ground truth teacher MLP with random weights, lmKANs achieved up to 6× fewer FLOPs at matched accuracy. On randomly displaced methane configurations, efficiency improved by up to 78×, or even more, for the Distances representation. Within convolutional networks, the gains were smaller but still significant: 1.6–2.1× on CIFAR-10 and 1.7× on ImageNet.
@@ 7 Acknowledgments
S.P. thanks Prashanth Kanduri, Nicholas J. Browning, and Henrique Mendonça for providing initial knowledge of CUDA when working on an unrelated project, Bob Crovella, whose instrumental lectures are available online [^4] https://www.youtube.com/playlist?list=PL6RdenZrxrw-zNX7uuGppWETdxt_JxdMj, and all the teachers of the CUDA course at Yandex School of Data Analysis.
@@ Appendix A Additional related work
The proposed lookup multivariate Kolmogorov-Arnold Networks fit within the general idea of conditional computation Bengio et al. [2013, 2015], which suggests using only part of a model’s parameters at inference time. The most popular subclass of such architectures is the Mixture-of-Experts (MoE) family Jacobs et al. [1991], Shazeer et al. [2017], Fedus et al. [2022], in which multiple experts, typically implemented as MLPs, process the current input only if they are deemed relevant.
In the proposed method, the spline lookup tables enable $\mathcal{O}(1)$ computations, which underpin the observed efficiency of our approach. For a function parametrized in this way, its value at any given point depends only on the small subset of parameters that control the function’s behaviour within the corresponding grid interval; all other parameters remain inactive, similarly to other conditional-computation methods.
@@
Outside the scope of general high-dimensional mappings, spline lookup tables have found several applications in machine learning. For instance, Graph Neural Networks (GNNs) applied to point clouds in Euclidean space Wang et al. [2018] often have trainable filters as part of the model. Such filters are one- or three-dimensional functions whose inputs are inter-point distances or displacement vectors. Sometimes these functions are splined Batatia et al. [2022] post-hoc, after the fitting of the model is finished.
@@ Appendix B More details on function parametrization
Finally, there are two linear basis functions on the left-most and right-most infinite intervals. We define them to be $0$-s in the left-most and right-most grid points, linearly increasing to $1$-s at the left and right ghost points and continuing to left and right infinities with the same slope, respectively.
@@
A reader may ask themselves a question of what is the difference between the proposed construction involving a function $f(x)$ defined on static percentile grid on $(-\infty,\infty)$ and a more simple approach which involves first mapping of an input to $(0,1)$ by $x^{\prime}=\sigma(x)$, and next computing a piecewise linear function $g(x^{\prime})$ defined on a uniform grid on $[0,1]$.
The answer is that the resulting functions in the initial space, $f(x)$ and $f^{\prime}(x)=g(\sigma(x))$, are not identical. Specifically, the $f^{\prime}(x)$ is no longer piecewise linear on each grid interval, with the most notable difference being the behavior when $x\rightarrow\pm\infty$. Asymptotes of $f(x)$ are linear, while asymptotes of $f^{\prime}(x)$ are constant.
The concern with horizontal asymptotes is that they can cause the vanishing gradient problem, which is believed to be one of the reasons why the ReLU Glorot et al. [2011] activations work better compared to the Tanh. Given these considerations, we chose the direct application of the static percentile grid for practical implementation.
@@ Appendix C More details on Hessian regularization
For the mixed derivative our finite-differences scheme is the following:
@@ E.1 Preconditioning
where the lmKAN weight, $\gamma$, is first set to 0 and then is gradually increased in our multistaged fitting procedure described later. In the case of ReLU-last preconditioning of Eq. 13a, we insert ReLU into all the layers except the last one; for ReLU-first preconditioning of Eq. 13b, we insert ReLU into all the layers except the first one. Therefore, at the beginning, when the lmKAN weight is zero, the model is equivalent to a pure MLP-based one for both types of preconditioning.
A merit of ReLU-first preconditioning is that during inference the whole Eq. 13b can be absorbed into a single lmKAN layer whenever the number of grid intervals $G$ is even, that is, when the origin is one of the grid points, see more details in appendix F.1. Thus, this type of preconditioning does not increase inference cost in any way. This is an advantage over the original KAN preconditioning scheme Liu et al. [2024], which requires the additional computation of the computationally expensive transcendental function SiLU for each edge at inference.
Because of the possibility of such an absorption, the total inference flops of a ReLU-first preconditioned lmKAN layer is 2x those of a linear layer of the same shape, while for the ReLU-last preconditioning, the slowdown factor is 3x, taking into account the linear branch.
@@ E.2 Fitting procedure
Phase II - turning on lmKAN: $\gamma$ is linearly (over time) increased from 0 to $0.3$. After that, there is some part with the constant $\gamma=0.3$. At this phase, we use very strong (= with very high coefficient $\lambda$) Hessian regularization introduced in Appendix. C. For all the subsequent phases, lmKAN weight $\gamma$ is fixed at $0.3$. The pipeline is also stable if increasing $\gamma$ to $1.0$, but in a few, though non-systematic, experiments, we found that using $0.3$ value leads to slightly better final accuracy.
@@
An example of the described fitting procedure for one of the training runs we did for numerical experiments described in Sec. 4.1 of the main text and in the Appendix. F.2 is given in Fig. 17.
@@ E.3 Comparison of the preconditioning schemes
We fitted lmKAN models with both types of preconditioning for the CIFAR-10 dataset and for general function approximation. The results are given in Fig. 18 and Fig. 19. Overall, the ReLU-last type of preconditioning appeared to lead to slightly more accurate models, but this small gain in accuracy doesn’t justify additional computational cost.
@@ F.1 General details about the benchmarking protocols
For lmKANs, when using ReLU-first preconditioning (see more details in Appendix. E), we absorb all the expression in Eq. 13b into the weights of the lmKAN layer. This modification requires to updated the lmKAN 2D functions as $f(x_{1},x_{2})\leftarrow\gamma f(x_{1},x_{2})+w_{1}\text{ReLU}(x_{1})+w_{2}\text{ReLU}(x_{2})$. Our construction allows for doing this absorption exactly whenever the origin is one of the grid points, which, in turn, is the case when the number of grid intervals $G$ is even. On top of that, batch normalizations are absorbed similarly to MLPs. We do not compile lmKAN models.
@@ F.2 General Function Approximation
There is a trend that the relative performance of lmKANs improves with the scale. It is clearly seen on the MSE vs FLOPs panel. On the MSE vs H100 wall clock time panel, it is first masked by the non-homogeneous efficiency of the code, but next still reveals itself for the largest hidden dimensions.
@@ F.3.1 The Distances Polynomials representation
The exact form of these polynomials is given in Eq. 14 and Eq. 15, where $x_{1}$, $x_{2}$, … $x_{10}$ correspond to interatomic distances $CH_{1}$, $CH_{2}$, $CH_{3}$, $CH_{4}$, $H_{1}H_{2}$, $H_{1}H_{3}$, $H_{1}H_{4}$, $H_{2}H_{3}$, $H_{2}H_{4}$, and $H_{3}H_{4}$ respectively. The presented polynomials form invariant generators Derksen and Kemper [2015] of the group corresponding to arbitrary permutations of the hydrogen atoms. Therefore, the Distances Polynomials representation preserves all the information about the initial $CH_{4}$ molecule.
@@ F.3.2 Ablations
The ablation study on the effect of Hessian regularization is given in Fig. 21, and the one of the effect of the number of grid intervals $G$ in Fig. 22.
@@ F.4 CIFAR-10
CIFAR-10 augmentation pipeline On top of these we use MixUP Zhang et al. [2017] ($\alpha=0.2$) and CutMix Yun et al. [2019] ($\beta=1.0$) augmentations, both with 50% probability. When fitting the families of convolutional neural networks described in the main text, we use the above pool of augmentations consistently for MLP-based and lmKAN-based CNNs. For all the details about the fitting procedures see the configurational files attached to these appendices.

### v2

@@ 4.3.1 CIFAR-10
Our backbone architecture consists of five $2\times 2$ convolutions, each with stride $2$, and two fully connected layers at the end. Since the resolution of CIFAR-10 images is $32\times 32$, where $32=2^{5}$, five $2\times 2$ convolutions with stride $2$ transform the spatial dimensions of an image exactly to $1\times 1$. All the layers use the same width (= number of filters in case of convolutions, and hidden dimension in case of fully connected layers), which we vary for both families of the models. In other aspects, the models are similar to those we employed in previous sections - we use batch normalizations with affine transforms for MLP-CNNs, and without for lmKAN-CNNs; MLP-CNNs use ReLU activations, while lmKAN-CNNs do not require additional activation layers.
The dataset comes with pre-defined full training and test subsets. We split the full training subset into training and validation parts in a 90%/10% ratio. Our augmentation pipeline consists of established techniques, such as RandAugment [53], MixUp [54], CutMix [55], and a few others. Further details about the augmentations and other aspects of our fitting setup are available in Appendix F.4.
Our findings are illustrated in the upper row of Fig. 11. Similarly to previous experiments, lmKAN-based CNNs were found to be more FLOPs efficient compared to classical MLP-based CNNs at the same accuracy level. The observed speed-up factor of 1.6-2.1× is not so dramatic as we report in sections 4.1 and 4.2, but still substantial. We have not yet implemented dedicated CUDA kernels for efficient inference of lmKAN-based convolutions. Any type of convolution can be cast to a fully connected layer by the corresponding memory manipulations, which we followed during fitting.
@@
Appendix F.4 contains ablation studies on 1) the effect of the number of grid intervals $G$ and 2) the effect of the strength of Hessian regularization $\lambda$.
@@ 4.3.2 ImageNet
To limit the computational cost, we downsampled the images to $81\times 81$ ($81=3^{4}$) pixels. Our backbone, illustrated in the right panel of Fig. 10, consists of four convolutional layers with the $3\times 3$ kernel size and stride $3$ and two fully connected layers. In contrast to the CIFAR-10 experiment, we progressively increase the number of filters as the spatial resolution of the image decreases through the neural network. Our augmentation pipeline consists of the same techniques as the ones we use for CIFAR-10, but with different hyperparameters; see more details in Appendix F.5. Since the test subset is not publicly available, we use the validation accuracy as the target metric. The performance of the models is summarized in the bottom row of Fig. 11. The observed efficiency gains of 1.7× are in line with those for CIFAR-10.
@@ 4.4 Comparison with FastKAN
We use the training script [^3] https://github.com/ZiyaoLi/fast-kan/blob/master/examples/train_cifar10.py for the CIFAR-10 dataset available in the FastKAN GitHub repository [56] as the basis for the comparison of lmKAN and FastKAN. However, we provide several modifications to the pipeline.
@@
We extend the script by the same augmentation pipeline as we used in Sec. 4.3 for the CIFAR-10 dataset. We observed that because of augmentations, one has to fit the model for a longer time, so the training budget was substantially increased. Additional changes are detailed in Appendix F.6.
Figure 12: Comparison of lmKAN and FastKAN within the fully connected framework on the CIFAR-10 dataset. Both models were fit with 1k and 10k epochs.
These modifications significantly improve the performance of FastKAN models ($54-55\%$ validation accuracy in the original script), see Fig. 12. Furthermore, even the MLP baseline of the same shape yields better accuracy compared to the performance of the FastKAN model in the original script.
@@
The first observation is that the performance of FastKAN models degrades for excessively fine grid resolutions. For the training budget of one thousand epochs, the final model is even less accurate than the MLP baseline. For the ten thousand epochs, the effect is less pronounced, but it still takes place. For lmKAN models, this degradation is much less severe, if present at all.
The number of parameters per function scales quadratically with grid resolution for lmKANs and linearly for FastKANs. Thus, although the rightmost lmKAN model in Fig. 12 uses more parameters per function, it operates on a much coarser grid — only $G=16$ grid intervals per dimension, compared with num_grids=200num\_grids=200 for FastKANs. This suggests that, for a comparable parameter budget per function, this coarser grid enforces a lower-frequency function class and, in turn, contributes to the superior training stability observed for lmKANs.
Another distinct feature of Fig. 12 is that lmKANs achieve notably better accuracy compared to FastKANs, even when the latter are evaluated with a very rich parametrization of inner functions. Additionally, note that the number of trainable functions $N_{in}/2]N_{out}$ and $N_{in}N_{out}$ in an lmKAN and FastKAN layer, respectively. Thus, the same number of parameters per function corresponds to a two times larger FastKAN model overall.
To conclude, these findings reinforce the intuitive considerations given in Sec. 3 and suggest that building blocks of multivariate trainable functions are indeed more effective.
@@ 6 Summary
High-dimensional linear mappings are the cornerstone of modern deep learning, dominating both the parameter count and the computational cost in most models. We introduce lookup multivariate Kolmogorov-Arnold Networks (lmKANs) as a drop-in replacement that offers a substantially better capacity–inference cost trade-off. Across all experiments, lmKANs were Pareto-optimal in the inference FLOPs–performance plane. The efficiency gains are task-dependent: for general high-dimensional function approximation, modeled as a distillation from a large ground-truth teacher MLP with random weights, lmKANs achieved up to 6× fewer FLOPs at matched accuracy. On randomly displaced methane configurations, efficiency improved by up to 78×, or even more, for the Distances representation. Within convolutional networks, the gains were smaller but still significant: 1.6–2.1× on CIFAR-10 and 1.7× on ImageNet.
@@ 7 Acknowledgments
S.P. thanks Prashanth Kanduri, Nicholas J. Browning, and Henrique Mendonça for fruitful discussions regarding CUDA, Bob Crovella, whose instrumental lectures are available online [^4] https://www.youtube.com/playlist?list=PL6RdenZrxrw-zNX7uuGppWETdxt_JxdMj, and all the teachers of the CUDA course at Yandex School of Data Analysis.
@@ Appendix A Additional related work
The proposed lookup multivariate Kolmogorov-Arnold Networks fit within the general idea of conditional computation [57, 58], which suggests using only part of a model’s parameters at inference time. The most popular subclass of such architectures is the Mixture-of-Experts (MoE) family [59, 60, 61], in which multiple experts, typically implemented as MLPs, process the current input only if they are deemed relevant.
In the proposed method, the spline lookup tables enable $\mathcal{O}(1)$ computations, which underpin the observed efficiency of our approach. For a function parametrized in this way, its value at any given point depends only on the small subset of parameters that control the function’s behavior within the corresponding grid interval; all other parameters remain inactive, similarly to other conditional-computation methods.
@@
Outside the scope of general high-dimensional mappings, spline lookup tables have found several applications in machine learning. For instance, Graph Neural Networks (GNNs) applied to point clouds in Euclidean space [62] often have trainable filters as part of the model. Such filters are one- or three-dimensional functions whose inputs are inter-point distances or displacement vectors. Sometimes these functions are splined [63] post-hoc, after the fitting of the model is finished.
@@ Appendix B More details on function parametrization
Finally, there are two linear basis functions on the left-most and right-most infinite intervals. We define them to be zeros in the left-most and right-most grid points, linearly increasing to ones at the left and right ghost points and continuing to left and right infinities with the same slope, respectively.
@@
A reader may ask themselves a question of what is the difference between the proposed construction involving a function $f(x)$ defined on static percentile grid on $(-\infty,\infty)$ and a more simple approach which involves first mapping of an input to $(0,1)$ by $x^{\prime}=\sigma(x)$, and then computing a piecewise linear function $g(x^{\prime})$ defined on a uniform grid on $[0,1]$.
The answer is that the resulting functions in the initial space, $f(x)$ and $f^{\prime}(x)=g(\sigma(x))$, are not identical. Specifically, $f^{\prime}(x)$ is no longer piecewise linear on each grid interval, with the most notable difference being the behavior when $x\rightarrow\pm\infty$. Asymptotes of $f(x)$ are linear, while asymptotes of $f^{\prime}(x)$ are constant.
The concern with horizontal asymptotes is that they can cause the vanishing gradient problem, which is believed to be one of the reasons why the ReLU [34] activations work better than tanh. Given these considerations, we chose the direct application of the static percentile grid for practical implementation.
@@ Appendix C More details on Hessian regularization
For the mixed derivative, our finite-differences scheme is the following:
@@ E.1 Preconditioning
where the lmKAN weight, $\gamma$, is initially set to 0 and then gradually increased in our multistaged fitting procedure described later. In the case of ReLU-last preconditioning of Eq. 12a, we insert ReLU into all the layers except the last one; for ReLU-first preconditioning of Eq. 12b, we insert ReLU into all the layers except the first one. Therefore, at the beginning, when the lmKAN weight is zero, the model is equivalent to a pure MLP-based one for both types of preconditioning.
A merit of ReLU-first preconditioning is that during inference the whole Eq. 12b can be absorbed into a single lmKAN layer whenever the number of grid intervals $G$ is even, that is, when the origin is one of the grid points, see more details in Appendix F.1. Thus, this type of preconditioning does not increase inference cost in any way. This is an advantage over the original KAN preconditioning scheme [11], which requires the additional computation of the computationally expensive transcendental function SiLU for each edge at inference.
Because of the possibility of such an absorption, the total inference FLOPs of a ReLU-first preconditioned lmKAN layer is 2× those of a linear layer of the same shape, while for the ReLU-last preconditioning, the slowdown factor is 3×, taking into account the linear branch.
@@ E.2 Fitting procedure
Phase II - turning on lmKAN: $\gamma$ is linearly (over time) increased from 0 to $0.3$. After that, there is some part with the constant $\gamma=0.3$. At this phase, we use very strong (= with a very high coefficient $\lambda$) Hessian regularization introduced in Appendix C. For all the subsequent phases, lmKAN weight $\gamma$ is fixed at $0.3$. The pipeline is also stable if $\gamma$ is increased to $1.0$, but in a few experiments we found that using $0.3$ value leads to slightly better final accuracy.
@@
An example of the described fitting procedure for one of the training runs we did for numerical experiments described in Sec. 4.1 of the main text and in the Appendix F.2 is given in Fig. 17.
@@ E.3 Comparison of the preconditioning schemes
We fitted lmKAN models with both types of preconditioning for the CIFAR-10 dataset and for general function approximation. The results are given in Fig. 18 and Fig. 19. Overall, the ReLU-last type of preconditioning appeared to lead to slightly more accurate models, but this small gain in accuracy does not justify additional computational cost.
@@ F.1 General details about the benchmarking protocols
For lmKANs, when using ReLU-first preconditioning (see more details in Appendix E), we absorb all the expression in Eq. 12b into the weights of the lmKAN layer. This modification requires updating the lmKAN 2D functions as $f(x_{1},x_{2})\leftarrow\gamma f(x_{1},x_{2})+w_{1}\text{ReLU}(x_{1})+w_{2}\text{ReLU}(x_{2})$. Our construction allows this absorption to be done exactly whenever the origin is one of the grid points, which, in turn, is the case when the number of grid intervals $G$ is even. On top of that, batch normalizations are absorbed similarly to MLPs. We do not compile lmKAN models.
@@ F.2 General Function Approximation
There is a trend that the relative performance of lmKANs improves with the scale. It is clearly seen on the MSE vs FLOPs panel. On the MSE vs H100 wall-clock time panel, it is first masked by the non-homogeneous efficiency of the code, but next still reveals itself for the largest hidden dimensions.
@@ F.3.1 The Distances Polynomials representation
The exact form of these polynomials is given in Eq. 13 and Eq. 14, where $x_{1}$, $x_{2}$, … $x_{10}$ correspond to interatomic distances $CH_{1}$, $CH_{2}$, $CH_{3}$, $CH_{4}$, $H_{1}H_{2}$, $H_{1}H_{3}$, $H_{1}H_{4}$, $H_{2}H_{3}$, $H_{2}H_{4}$, and $H_{3}H_{4}$ respectively. The presented polynomials form invariant generators [65] of the group corresponding to arbitrary permutations of the hydrogen atoms. Therefore, the Distances Polynomials representation preserves all the information about the initial $CH_{4}$ molecule.
@@ F.3.2 Ablations
The ablation study on the effect of Hessian regularization is given in Fig. 21, and the effect of the number of grid intervals $G$ in Fig. 22.
@@ F.4 CIFAR-10
CIFAR-10 augmentation pipeline On top of these, we use MixUp [54] ($\alpha=0.2$) and CutMix [55] ($\beta=1.0$) augmentations, both with 50% probability. When fitting the families of convolutional neural networks described in the main text, we use the above pool of augmentations consistently for MLP-based and lmKAN-based CNNs. For all the details about the fitting procedures see the configuration files attached to these appendices.